# Lightning-Fast Image Inversion and Editing for Text-to-Image Diffusion Models

### v4

@@ 5.3 Real-Time Image Editing
We now evaluate the effect of inverting images with GNRI, compared to SoTA inversion baselines. We show that GNRI outperforms all baselines in editing performance while requiring the shortest time. This opens the door for real-time editing capabilities. We present below both qualitative and quantitative results.
Qualitative results. Figure 5(left) gives a qualitative comparison between SoTA approaches and GNRI in the context of real image editing. GNRI excels in accurately editing with target prompts producing images with both high fidelity to the input image and adherence to the target prompt. The examples illustrate how alternative approaches may struggle to retain the original structure or tend to overlook crucial components specified in the target prompt. As an example, in the first row of Figure 5(left), GNRI exclusively converts the cat into a dog without changing the text in the sign whereas other methods either fail to generate a proper dog or change the text in the sign. Results for SD2.1 in Appendix. G.
User Study. We further evaluated editing quality using human raters. We followed the evaluation process of Mokady et al. (2023) and asked human raters to rank edits made by the three leading methods: Ours, ReNoise (Garibi et al., 2024), and ExactDPM (Hong et al., 2024b). Unfortunately, the code for TurboEdit (Wu et al., 2024) was not available at the time of our evaluation. 100 images were rated, 12 images were provided by the authors of (Mokady et al., 2023) and the rest were randomly selected from COCO (Lin et al., 2014) and PieBench (Ju et al., 2023) benchmarks. Three raters for each image were recruited through Amazon Mechanical Turk and were tasked to choose the image that better applies the requested text edit while preserving most of the original image. GNRI was preferred with 71.6% preference compared to ReNoise’s (Garibi et al., 2024) 15.9% and ExactDPM (Hong et al., 2024b) 12.5%.
Quantitative results. Following (Wallace et al., 2022; Huberman-Spiegelglas et al., 2023; Garibi et al., 2024), we evaluate the results using two complementary metrics: LPIPS (Zhang et al., 2018) to quantify the extent of structure preservation (lower is better) and a CLIP-based score to quantify how well the generated images comply with the text prompt (higher is better). Metrics are averaged across 100 MS-COCO images. Figure 5(right) illustrates that editing with GNRI yields a superior CLIP and LPIPS score, demonstrating the ability to perform state-of-the-art real image editing with superior fidelity. Following (Wu et al., 2024) we also evaluated our approach on the newly introduced PieBench (Ju et al., 2023) dataset. Table 1 shows that our method more effectively follows the text guidance while preserving the background, outperforming current SoTA techniques. These evaluations further affirm the findings derived from the user study.
@@
ExactDPM (Hong et al., 2024b) 24.54 59.88 36.49 69.18 23.77 21.23 15.11 0.445
@@ 5.4 Seed Interpolation and Rare Concept Generation
We evaluate the benefit of our inversion in a second task, of selecting latent seeds for image generation. It has been shown (Samuel et al., 2024; 2023) that text-to-image diffusion models poorly generated rare concepts, and this can be improved by selecting better noise latent $z_{T}$. To address this issue, SeedSelect (Samuel et al., 2024) takes a few images of a rare concept as input and uses a diffusion inversion module to iteratively refine the obtained seeds. These refined seeds are then used to generate new plausible images of the rare concept. NAO (Samuel et al., 2023) extends this by introducing new paths and centroids for seed initialization. Both methods rely on DDIM Inversions, crucial for initial seed evaluation. GNRI provides an alternative for precise inversion seeds, aiming for improved image quality and semantic accuracy. It’s important to highlight that most inversion techniques (Mokady et al., 2023; Wallace et al., 2022; Huberman-Spiegelglas et al., 2023; Brack et al., 2024) are not applicable in this context as they necessitate extra parameters for image reconstruction, which impedes the straightforward implementation of interpolation.
Interpolation and centroid finding: We evaluate GNRI by following the experimental protocol of (Samuel et al., 2023) and compare images generated by four methods: DDIM, ReNoise (Garibi et al., 2024), ExactDOM (Hong et al., 2024a) and GNRI. Evaluation is conducted based on FID score and image accuracy, assessed using a pre-trained classifier. See details in (Samuel et al., 2023). Results are presented in Table 2. Notably, initializing with GNRI seeds consistently results in higher-quality images both in interpolation paths and seed centroids.
@@ 6 Summary
Image inversion in diffusion models is vital for various applications like image editing, semantic augmentation, and generating rare concept images. Current methods often sacrifice inversion quality for computational efficiency, requiring significantly more compute resources for high-quality results. This paper presents Guided Newton-Raphson Inversion (GNRI), a novel iterative approach that balances rapid convergence with superior accuracy, execution time, and memory efficiency. Using GNRI opens the door for high-quality real-time image editing due to its efficiency and speed. GNRI requires no model training, fine-tuning, prompt optimization, or additional parameters, and is compatible with all pre-trained diffusion models and deterministic schedulers.
@@ Appendix A Newton method for Multivariable Scalar Function
When $f>0$ and does not have zero crossings, the NR procedure may converge to a (non-zero) minima. A common example is the function $f(x)=x^{2}+1$ that has no real roots, but NR converges to its minimum. We believe that an analysis of this type of convergence within the Newton-Raphson scheme, falls outside the scope of this paper.
Appendix B GNRI Scheme Contraction Mapping
@@
$z^{k+1}_{i}=z^{k}_{i}-\frac{1}{D}\leavevmode\nobreak\ \frac{{\mathcal{F}}(z^{k})}{g_{i}(z^{k})}$ (B1)
@@
$\tilde{f}_{i}(z^{k}):=z^{k}_{i}-\frac{1}{D}\leavevmode\nobreak\ \frac{{\mathcal{F}}(z^{k})}{g_{i}(z^{k})}$ (B2)
@@
$\tilde{r}(z^{k+1})$ $=||\tilde{f}(z^{k+1})-z^{k+1})||=||f(z^{k+1})-\tilde{f}(z^{k})||$ (B4)
@@
$\tilde{r}(z^{k+1})\leq\rho\leavevmode\nobreak\ \tilde{r}(z^{k})$ (B5)
and the contraction mapping is satisfied if $\tilde{r}$ follows an exponential decay ($r(z^{k})\leq\rho^{k}\leavevmode\nobreak\ r(z^{0})$), with some $\rho<1$. Note that $\tilde{r}(z^{k})$ given in Eq. B3 can be computed at each iteration step. In Fig. B1 we plot $\tilde{r}(z^{k})$, tested on 5,000 image caption pairs from the COCO dataset (the same as used in Sec. 5.1) showing that $\tilde{r}(z^{k})$ indeed decays exponentially (linearly in log-scale), till convergence. This behaviour therefore implies that GNRI iteration is a contraction mapping. Note that our contraction mapping does not converge to zero, namely the fixed point, since our objective function $\mathcal{F}$ may not have an exact solution. Computing contraction rate $\rho$ for GNRI shows $\rho\approx 0.50$ (at the first iteration) while $\rho\approx 0.90$ for FPI indicating the faster convergence of GNRI.
@@
Figure B1: Contraction Mapping. Contraction mapping for two fast diffusion models are shown (in log-scale), at each time-step. The scheme converges in 2-3 iterations with a notable contraction occurring in the first iteration, assisting the rapid convergence. $t=0$ is close to the image side and $t=3$ close to seed.
Appendix C Failure Cases
@@
Here we provide information about an analysis to find failure cases for convergence. We ran inversion and reconstruction, in scale, on COCO2017 (118k caption-image pairs) and found that in 95.4% of cases, GNRI successfully converged to a solution. Specifically, residuals went down from $\sim 1$ (for DDIM) to $<10^{-4}$ indicating convergence, and the PSNR was $>25.7$ indicating good solutions. The remaining $4.6\%$ were samples with incorrect captions, see Fig. SC1. These results imply that lack of convergence may indicate text and image miss-alignment which we consider for future work.
@@
Figure C1: Failure cases: GNRI fails to converge where the prompt and image do not align.
@@
Figure C2: The influence of $\lambda$ on reconstruction performance. We observe that employing no regularization ($\lambda=0$) leads to poor reconstruction, while $\lambda=0.1$ typically yields the highest reconstruction accuracy. As $\lambda$ increases, reconstruction accuracy declines, possibly because assigning excessive weight to the prior undermines the root-finding objective.
Appendix D Prior trade-off parameter
@@
Figure C2 shows the impact of the weight parameter $\lambda$ on the reconstruction accuracy. We observe that using no regularization ($\lambda=0$) results in poor reconstruction, whereas setting $\lambda=0.1$ achieves the highest reconstruction accuracy, underscoring the importance of the prior in the objective function. There is a gradual decrease with larger $\lambda$ values, where at $\lambda=1$ the PSNR starts to decline faster, likely because placing too much emphasis on the prior compromises the residual root-finding objective.
@@
Figure D1: (Left) Inversion Results: Mean reconstruction quality (y-axis, PSNR) and runtime (x-axis, seconds) on the COCO2017 validation set. Our method achieves high PSNR while reducing inversion-reconstruction time by a factor of ×$\times 4$ to ×$\times 14$.(Right) Evaluation of editing performance: GNRI achieves superior CLIP and LPIPS scores, indicating better compliance with text prompts and higher structure preservation.
Appendix E Seed Exploration & Rare concept generation
@@
Figure E1: Generating rare concepts based on a few examples with the method of (Samuel et al., 2023) that heavily depends on the diffusion-inversion quality. In our comparison, we evaluate GNRI alongside ReNoise (refer to the discussion in the main paper). ReNoise frequently struggles to produce a realistic representation of certain objects (such as Oxygen-mask or Patas Monkey) or an accurate depiction of specific concepts (like Tiger-cat or Pay-phone). However, the use of GNRI rectifies these issues in the results. For a detailed quantitative comparison, please refer to the main paper. See the main paper for a quantitative comparison.
@@
Figure SE1 further displays results for the rare-concept generation task introduced in (Samuel et al., 2024; 2023). The objective is to enable the diffusion model to generate concepts rarely seen in its training set by using a few images of that concept. Both methods in (Samuel et al., 2024; 2023) utilize diffusion inversion for this purpose. In the main paper, we presented experiments showcasing the impact of our new inversion process on the outcomes of (Samuel et al., 2024; 2023).
@@
Figure E2: Qualitative comparison for image editing: Illustrates how alternative approaches may struggle to preserve the original structure or overlook crucial components specified in the target prompt, while GNRI succeeds in editing the image properly.
@@
Figure E3: Various editing with same input: Note the GNRI capability in both subtle and extensive changes as one would expect from the particular prompt change.
@@
Appendix F Additional Quantitative Results
@@
In this section, we present a quantitative comparison between GNRI and baseline methods, utilizing latent diffusion SD2.1 (Rombach et al., 2022). Figure D1(a) illustrates the image inversion-reconstruction performance, while Figure D1(b) highlights the editing performance, evaluated using LPIPS and CLIP-score metrics. The results demonstrate that GNRI surpasses all other inversion methods in terms of reconstruction quality, which subsequently improves editing performance.
Appendix G Additional Qualitative Results
@@
In this section, we present additional qualitative comparisons involving GNRI and baseline methods, using SD2.1.
@@
Figures E2 provide further comparisons for real-image editing. These figures illustrate how alternative approaches may struggle to preserve the original structure or overlook crucial components specified in the target prompt. For example, n the first row of Figure E2, GNRI correctly transforms the pizza into bread, while other methods generate a bad looking bread. Row three shows an example of GNRI success (bananas to oranges), where other alternatives totally fail.
@@
Figure E3 shows the outcomes of various edits on the same image, providing additional confirmation that inversion with GNRI yields modification of the pertinent parts in the image while maintaining the original structure.

### v5

@@ 5.3 Real-Time Image Editing
We now evaluate the effect of inverting images with GNRI, compared to SoTA inversion baselines. We show that GNRI outperforms all baselines in editing performance while requiring the shortest time. This opens the door for real-time editing capabilities. We present below both qualitative and quantitative results. For Neural Function Evaluations (NFE) comparison see Appendix H.
Qualitative results. Figure 5(left) gives a qualitative comparison between SoTA approaches and GNRI in the context of real image editing. GNRI excels in accurately editing with target prompts producing images with both high fidelity to the input image and adherence to the target prompt. The examples illustrate how alternative approaches may struggle to retain the original structure or tend to overlook crucial components specified in the target prompt. As an example, in the first row of Figure 5(left), GNRI exclusively converts the cat into a dog without changing the text in the sign whereas other methods either fail to generate a proper dog or change the text in the sign. Results for SD2.1 in Appendix H. For additional qualitative results see also Appendix H,G and K. User Study. We further evaluated editing quality using human raters. We followed the evaluation process of Mokady et al. (2023) and asked human raters to rank edits made by the three leading methods: Ours, ReNoise (Garibi et al., 2024), and ExactDPM (Hong et al., 2024). Unfortunately, the code for TurboEdit (Wu et al., 2024) was not available at the time of our evaluation. 100 images were rated, 12 images were provided by the authors of (Mokady et al., 2023) and the rest were randomly selected from COCO (Lin et al., 2014) and PieBench (Ju et al., 2023) benchmarks. Three raters for each image were recruited through Amazon Mechanical Turk and were tasked to choose the image that better applies the requested text edit while preserving most of the original image. GNRI was preferred with 71.6% preference compared to ReNoise’s (Garibi et al., 2024) 15.9% and ExactDPM (Hong et al., 2024) 12.5%.
Quantitative results. Following (Wallace et al., 2023; Huberman et al., 2023; Garibi et al., 2024), we evaluate the results using two complementary metrics: LPIPS (Zhang et al., 2018) to quantify the extent of structure preservation (lower is better) and a CLIP-based score to quantify how well the generated images comply with the text prompt (higher is better). Metrics are averaged across 100 MS-COCO images. Figure 5(right) illustrates that editing with GNRI yields a superior CLIP and LPIPS score, demonstrating the ability to perform state-of-the-art real image editing with superior fidelity. Following (Wu et al., 2024) we also evaluated our approach on the newly introduced PieBench (Ju et al., 2023) dataset. Table 1 shows that our method more effectively follows the text guidance while preserving the background, outperforming current SoTA techniques. These evaluations further affirm the findings derived from the user study.
@@
ExactDPM (Hong et al., 2024) 24.54 59.88 36.49 69.18 23.77 21.23 15.11 0.445
EditFriendly (Huberman et al., 2023) 24.55 91.88 95.58 81.57 23.97 21.03 32.3 23.1
@@ 5.4 Seed Interpolation and Rare Concept Generation
We evaluate the benefit of our inversion in a second task, of selecting latent seeds for image generation. It has been shown (Samuel et al., 2024; 2023) that text-to-image diffusion models poorly generated rare concepts, and this can be improved by selecting better noise latent $z_{T}$. To address this issue, SeedSelect (Samuel et al., 2024) takes a few images of a rare concept as input and uses a diffusion inversion module to iteratively refine the obtained seeds. These refined seeds are then used to generate new plausible images of the rare concept. NAO (Samuel et al., 2023) extends this by introducing new paths and centroids for seed initialization. Both methods rely on DDIM Inversions, crucial for initial seed evaluation. GNRI provides an alternative for precise inversion seeds, aiming for improved image quality and semantic accuracy. It’s important to highlight that most inversion techniques (Mokady et al., 2023; Wallace et al., 2023; Huberman et al., 2023; Brack et al., 2024) are not applicable in this context as they necessitate extra parameters for image reconstruction, which impedes the straightforward implementation of interpolation.
Interpolation and centroid finding: We evaluate GNRI by following the experimental protocol of (Samuel et al., 2023) and compare images generated by four methods: DDIM, ReNoise (Garibi et al., 2024), ExactDOM (Hong et al., 2024) and GNRI. Evaluation is conducted based on FID score and image accuracy, assessed using a pre-trained classifier. See details in (Samuel et al., 2023). Results are presented in Table 2. Notably, initializing with GNRI seeds consistently results in higher-quality images both in interpolation paths and seed centroids.
@@ 6 Summary
Image inversion in diffusion models is vital for various applications like image editing, semantic augmentation, and generating rare concept images. Current methods often sacrifice inversion quality for computational efficiency, requiring significantly more compute resources for high-quality results. This paper presents Guided Newton-Raphson Inversion (GNRI), a novel iterative approach that balances rapid convergence with superior accuracy, execution time, and memory efficiency. Using GNRI opens the door for high-quality real-time image editing due to its efficiency and speed. GNRI requires no model training, fine-tuning, prompt optimization, or additional parameters, and is compatible with all pre-trained diffusion and flow-matching models with deterministic schedulers.
Acknowledgments
We would like to express our sincere gratitude to Prof. Nir Sochen for the intriguing discussion and valuable insights.
@@ Appendix A Newton method for Multivariable Scalar Function
Appendix B Discussion on the Newton-Raphson method for functions without zero crossings
Note that our objective function in Eq. 9 is a non-negative function. When the target function $\mathcal{F}\geq 0$ and does not have zero crossings,
the NR procedure may converge to a (non-zero) minima. A common example is the function $f(x)=x^{2}+1$ that has no real roots, but NR converges to its minimum. Note that while the Newton-Raphson method is typically used to find zero crossings of a function, our approach applies it to drive convergence toward local minima of the objective function (being locally convex around this point). The analysis of this type of convergence within the Newton-Raphson framework is left for future numerical analysis research.
Appendix C GNRI Scheme Contraction Mapping
@@
$z^{k+1}_{i}=z^{k}_{i}-\frac{1}{D}\leavevmode\nobreak\ \frac{{\mathcal{F}}(z^{k})}{g_{i}(z^{k})}$ (C1)
@@
$\tilde{f}_{i}(z^{k}):=z^{k}_{i}-\frac{1}{D}\leavevmode\nobreak\ \frac{{\mathcal{F}}(z^{k})}{g_{i}(z^{k})}$ (C2)
@@
$\tilde{r}(z^{k+1})$ $=||\tilde{f}(z^{k+1})-z^{k+1})||=||f(z^{k+1})-\tilde{f}(z^{k})||$ (C4)
@@
$\tilde{r}(z^{k+1})\leq\rho\leavevmode\nobreak\ \tilde{r}(z^{k})$ (C5)
and the contraction mapping is satisfied if $\tilde{r}$ follows an exponential decay ($r(z^{k})\leq\rho^{k}\leavevmode\nobreak\ r(z^{0})$), with some $\rho<1$. Note that $\tilde{r}(z^{k})$ given in Eq. C3 can be computed at each iteration step. In Fig. C1 we plot $\tilde{r}(z^{k})$, tested on 5,000 image caption pairs from the COCO dataset (the same as used in Sec. 5.1) showing that $\tilde{r}(z^{k})$ indeed decays exponentially (linearly in log-scale), till convergence. This behaviour therefore implies that GNRI iteration is a contraction mapping. Note that our contraction mapping does not converge to zero, namely the fixed point, since our objective function $\mathcal{F}$ may not have an exact solution. Computing contraction rate $\rho$ for GNRI shows $\rho\approx 0.50$ (at the first iteration) while $\rho\approx 0.90$ for FPI indicating the faster convergence of GNRI.
@@
Figure C1: Contraction Mapping. Contraction mapping for two fast diffusion models are shown (in log-scale), at each time-step. The scheme converges in 2-3 iterations with a notable contraction occurring in the first iteration, assisting the rapid convergence. $t=0$ is close to the image side and $t=3$ close to seed.
Appendix D Failure Cases
@@
Here we provide information about an analysis to find failure cases for convergence. We ran inversion and reconstruction, in scale, on COCO2017 (118k caption-image pairs) and found that in 95.4% of cases, GNRI successfully converged to a solution. Specifically, residuals went down from $\sim 1$ (for DDIM) to $<10^{-4}$ indicating convergence, and the PSNR was $>25.7$ indicating good solutions. The remaining $4.6\%$ were samples with incorrect captions, see Fig. SD1. These results imply that lack of convergence may indicate text and image miss-alignment which we consider for future work.
@@
Figure D1: Failure cases: GNRI fails to converge where the prompt and image do not align.
@@
Figure D2: The influence of $\lambda$ on reconstruction performance. We observe that employing no regularization ($\lambda=0$) leads to poor reconstruction, while $\lambda=0.1$ typically yields the highest reconstruction accuracy. As $\lambda$ increases, reconstruction accuracy declines, possibly because assigning excessive weight to the prior undermines the root-finding objective.
Appendix E Prior trade-off parameter
@@
Figure D2 shows the impact of the weight parameter $\lambda$ on the reconstruction accuracy. We observe that using no regularization ($\lambda=0$) results in poor reconstruction, whereas setting $\lambda=0.1$ achieves the highest reconstruction accuracy, underscoring the importance of the prior in the objective function. There is a gradual decrease with larger $\lambda$ values, where at $\lambda=1$ the PSNR starts to decline faster, likely because placing too much emphasis on the prior compromises the residual root-finding objective.
@@
Figure E1: (Left) Inversion Results: Mean reconstruction quality (y-axis, PSNR) and runtime (x-axis, seconds) on the COCO2017 validation set. Our method achieves high PSNR while reducing inversion-reconstruction time by a factor of ×$\times 4$ to ×$\times 14$.(Right) Evaluation of editing performance: GNRI achieves superior CLIP and LPIPS scores, indicating better compliance with text prompts and higher structure preservation.
Appendix F Seed Exploration & Rare concept generation
@@
Figure F1: Generating rare concepts based on a few examples with the method of (Samuel et al., 2023) that heavily depends on the diffusion-inversion quality. In our comparison, we evaluate GNRI alongside ReNoise (refer to the discussion in the main paper). ReNoise frequently struggles to produce a realistic representation of certain objects (such as Oxygen-mask or Patas Monkey) or an accurate depiction of specific concepts (like Tiger-cat or Pay-phone). However, the use of GNRI rectifies these issues in the results. For a detailed quantitative comparison, please refer to the main paper. See the main paper for a quantitative comparison.
@@
Figure SF1 further displays results for the rare-concept generation task introduced in (Samuel et al., 2024; 2023). The objective is to enable the diffusion model to generate concepts rarely seen in its training set by using a few images of that concept. Both methods in (Samuel et al., 2024; 2023) utilize diffusion inversion for this purpose. In the main paper, we presented experiments showcasing the impact of our new inversion process on the outcomes of (Samuel et al., 2024; 2023).
@@
Figure F2: Qualitative comparison for image editing on the PIE-benchmark. Illustrates how alternative approaches may struggle to preserve the original structure or fail to edit the image, while GNRI succeeds in editing the image properly.
Figure F3: Qualitative comparison for image editing: Illustrates how alternative approaches may struggle to preserve the original structure or overlook crucial components specified in the target prompt, while GNRI succeeds in editing the image properly.
@@
Figure F4: Various editing with same input: Note the GNRI capability in both subtle and extensive changes as one would expect from the particular prompt change.
@@
Appendix G Additional Qualitative Results on PIE-Benchmark
Additional quantitative results on the PIE-Benchmark.
Figure F2 provides quantitative results on the PIE-Benchmark using Flux.1. It highlights how alternative methods often fail to maintain the original structure or miss key elements specified in the target prompt, whereas GNRI effectively achieves accurate and proper image editing.
Quantitative results with Stable Diffusion 2.1.
@@ Appendix G Additional Qualitative Results on PIE-Benchmark
We present a quantitative comparison between GNRI and baseline methods, utilizing latent diffusion SD2.1 (Rombach et al., 2022). Figure E1(a) illustrates the image inversion-reconstruction performance, while Figure E1(b) highlights the editing performance, evaluated using LPIPS and CLIP-score metrics. The results demonstrate that GNRI surpasses all other inversion methods in terms of reconstruction quality, which subsequently improves editing performance.
Appendix H Additional Qualitative and Quantitative Results -Comparison to Baselines
@@
In this section, we present additional qualitative comparisons involving GNRI and baseline methods.
@@
Figures F3 provide further comparisons for real-image editing, using SD2.1. These figures illustrate how alternative approaches may struggle to preserve the original structure or overlook crucial components specified in the target prompt. For example, n the first row of Figure F3, GNRI correctly transforms the pizza into bread, while other methods generate a bad looking bread. Row three shows an example of GNRI success (bananas to oranges), where other alternatives totally fail.
@@
Figure F4 shows the outcomes of various edits on the same image, providing additional confirmation that inversion with GNRI yields modification of the pertinent parts in the image while maintaining the original structure.
Table H1: PSNR vs. NFE results using SDXL Turbo on the COCO validation set. Our method achieves the highest PSNR with the fewest NFEs, compared to other methods.
PSNR NFE
ReNoise Garibi et al. (2024) 18.1 36
ExactDPM Hong et al. (2024) 20.0 40
TurboEdit Wu et al. (2024) 22 8
GNRI (Ours) 23.9 6
Table H2: Inversion comparison with Edit Friendly on SDXL Turbo (COCO validation set). While DDPM achieves higher PSNR through stochastic inversion, our deterministic approach (GNRI) provides competitive results with significantly faster inversion time (0.521s vs 32.3s).
PSNR Inv Time
EditFriendly Huberman et al. (2023) 26.1 32.3
GNRI (Ours) 23.9 0.521
Table H1 provides PSNR vs NFE (Neural Function Evaluations) results using SDXL Turbo on the COCO validation set. Our method requires 6 NFEs, calculated as 4 steps with an average of 1.5 NR iterations per step (some steps needing 1 iteration, others 2). As shown in the table below, our approach achieves the highest PSNR while using the fewest NFEs compared to other methods.
Table H2 compares inversion performance with Edit Friendly Huberman et al. (2023) using SDXL Turbo on the COCO validation set. Edit Friendly achieves the maximum PSNR possible given the VAE distortion, because it does a stochastic DDPM inversion, saving additional noise matrices at every step and “overfitting” the inverted image. As a result, it requires extensive inversion time (more than 30 seconds), whereas our GNRI reaches a relatively high PSNR in under a second. Our approach solves deterministic inversion (DDIM or Euler) making it suitable for applications like interpolation and rare concept generation, where DDPM is not feasible.
Appendix I Comparison to Anderson Acceleration
The guiding term introduced in Sec 4.2 might appear similar to Anderson acceleration used in Pan et al. (2023), which aims to allow more gradual changes from previous estimates. However, One key difference is that our guidance term introduced in Eq.(9) does not pull $z_{t}$ toward $z_{t-1}$ but towards the mean $\mu_{t}$. By that, it directs the Newton-Raphson method toward solutions having a high likelihood, using “side-information” about the noise. In contrast, Anderson acceleration relies on previous values of the implicit function $f(z_{i})$ and residuals but lacks awareness of the latent distribution (see for example, line 13 in Algorithm 1 of (Pan et al., 2023)).
Table I1: Comparison of Anderson Acceleration and the Guidance term for out-of-distribution COCO samples. A higher likelihood is better.
mean likelihood q($z_{t}|z_{t-1}$) PSNR Inv Time
Anderson Acceleration 0.34 18.3 Inf
Guidance Regularizer (Ours) 0.82 25.8 0.4 s
In GNRI, keeping $z_{t}$ close to $\mu$, encourages solutions that are more “in-distribution” and aligned with the diffusion model. In other words, if $z_{t-1}$ starts far out-of-distribution, keeping $z_{t}$ close to $z_{t-1}$ is likely to yield worse solutions compared to guiding $z_{t}$ toward inputs that the model expects to receive.
We illustrate this effect in table I1. We analyzed the inversion of 300 COCO images that the model would consider to be non-typical by selecting the samples with the lowest likelihood under $q(z_{0})$ (Eq.(1)). The table shows that GNRI drives the latents toward significantly higher likelihood solutions (mean $q(z_{t}|z_{t-1})$), leading to substantially better reconstructions (higher PSNR) than Anderson acceleration.
More broadly, our guidance term could potentially be applied to other methods, such as fixed-point iteration approaches like those proposed by Pan et al. It would also be interesting to find methods that can include more domain knowledge into the inversion process.
Appendix J Memory Usage
We compared the memory usage of our approach, with AIDI Pan et al. (2023) and ExactDPM (Zhang et al. (2023)) on a A100 GPU (40 GB VRAM) and a RTX 3090Ti GPU (24GB VRAM). Results are given in table J1. Our approach consumes about the same GPU memory as ExactDPM, however, it converges faster about x43 times than ExactDPM on an A100 GPU, and x25 faster on an RTX 3090 GPU.
Table J1: Memory usage and convergence time for different methods on A100 and RTX 3090Ti.
A100 (VRAM 40GB) RTX 3090Ti (VRAM 24GB)
Memory Inv Time Memory Inv Time
AIDI Pan et al. (2023) 15GB 5 sec 16 GB 13 sec
ExactDPM Hong et al. (2024) 30GB 17 sec 20 GB 25 sec
GNRI (Ours) 28GB 0.4 sec 19 GB 1 sec
Appendix K Multi Object Editing
We conducted several qualitative experiments to evaluate the capabilities of our method in editing multiple objects simultaneously. Our results demonstrate that the method can effectively handle complex scenes, allowing for the independent manipulation of several objects without compromising the quality of individual edits. As illustrated in Figure K1, our GNRI model successfully performs multi-object transformations while maintaining scene coherence and natural composition - from converting multiple dogs to cats while preserving the scene layout, to simultaneously transforming different food items (steak to apple, eggs to pancakes) on a plate. These results underscore GNRI’s robustness in handling multiple concurrent edits while maintaining high visual quality and semantic coherence across the transformed scenes.
Figure K1: Multi-object editing capabilities of GNRI. Top: Simultaneous transformation of multiple dogs into cats. Bottom: Concurrent food item conversion on a plate (steak→apple, eggs→pancakes), demonstrating preservation of scene composition and context.